solve using algebraic method. if possible check your solution
2x-5y=3
x-2y=7

Found 2 solutions by stanbon, macston:
Answer by stanbon(75887)

2x-5y=3
x-2y=7
---------
Multiply thru the bottom equation by 2 to get:
2x - 5y = 3
2x - 4y = 14
-------------
Subtract and solve for "y"::
y = 11
----
Solve for "x"::
x - 2y = 7
x - 2*11 = 7
x = 7 + 22 = 29
-------------------
Cheers,
Stan H.
-------------

Answer by macston(5194)

.
Second equation:
x-2y=7
x=2y+7
.
First equation:
2x-5y=3 Substitute for x.
2(2y+7)-5y=3
4y+14-53y=3
-y+14=3
-y=-11
y=11
.
x=2y+7
x=2(11)+7=29
.
ANSWER: x=29 and y=11
.
CHECK:
x-2y=7
29-2(11)=7
29-22=7
7=7
.
CHECK 2:
2x-5y=3
2(29)-5(11)=3
58-55=3
3=3